A bullet is fired through a board 10cm thick in such a way that the bullets line of motion is perpendicular to the face of the board. If the initial speed of the bullet is 400m/s and it emerges from the other side of the board with a speed of 200m/s, find the acceleration of the bullet as it passes through the board. Round two decimal places.


1
Expert's answer
2021-02-11T17:10:48-0500
"v_f^2-v_i^2=2as\\200^2-400^2=2(0.1)a\\a=-600000\frac{m}{s^2}"